# LetsGo: Large-Scale Garage Modeling and Rendering via LiDAR-Assisted Gaussian Primitives

### v1

@@ 7 Limitations and Discussions
Given that the large texture-less regions and repetitive patterns in garage scenes hinder the 3D reconstruction of expansive garages from images by SfM and MVS algorithms, this paper designs a Polar device using which the point cloud data of the large-scale garages can be easily captured and calibrated, and demonstrated that the collected LiDAR data successfully assist a suit of 3D Gaussian splatting algorithms, enabling high-quality rendering. Moreover, our lightweight renderer, combined with LOD techniques, facilitates real-time rendering of expansive garage environments on lightweight platforms. Although our rendering results exhibit commendable realism, our workflow still possesses certain limitations. Here we present a detailed analysis and explore further potential applications.
Firstly, our method focuses on large-scale garage scenes and relies on our lightweight 3D scanner to provide high-quality LiDAR point clouds, color images, and corresponding camera information. Although we have also validated the effectiveness of our method on other open-source datasets, it is important to explore the use of even lighter devices for scanning and rendering large-scale scenes, such as smartphones equipped with depth sensors. Additionally, we have currently collected data from five large-scale garages, but it is necessary to gather more garage data to contribute to the community and facilitate further research on garage modeling and rendering.
As a method based on image rendering, our approach achieves highly realistic rendering effects, almost indistinguishable from real scenes. However, the existing pipeline does not support modifying lighting conditions. This requires us to carefully design the shooting process according to the lighting conditions, limiting the applicability of our method in various scenarios. Recently, Gaussian-based work SuGaR [20] has introduced editing operations such as adjusting the size, position, and orientation of objects, but it still does not support modifying lighting information. Enabling rendering and editing of large-scale scenes under different lighting conditions is a meaningful direction that deserves further investigation. Our open-source dataset provides a foundation for the community to conduct research in these directions, allowing for advancements in this field.
We have demonstrated the applications of our method in automatic driving data generation, localization and navigation on lightweight devices, and the production of visual effects in films. These applications illustrate the excellent performance of our method in rendering quality. Furthermore, there are additional directions and applications worth exploring. Inspired by SMEREF [41], one potential research direction is to investigate streaming transmission methods for Gaussian kernels, enabling the distribution and on-the-fly rendering of large-scale scene data. Additionally, there are city generation methods such as InfiniCity [34] and CityDreamer [73], which leverage Generative Adversarial Networks (GANs) [19] to achieve rapid modeling of large-scale scenes. We intend to study large-scale scene generation based on 3D Gaussian representations. This fully explicit representation can be easily integrated into existing computer graphics workflows and achieve superior rendering effects.
@@ 8 Conclusion
This paper has contributed a handheld Polar device for data collection, a GarageWorld dataset, LiDAR-assisted Gaussian primitives for scene representation, and a lightweight rendering technique that allows web-based rendering on consumer-level devices. Benefiting from these innovations, we successfully reconstruct various garages with diverse and challenging environments, allowing real-time lightweight rendering from any viewpoint. Experimental results on the collected and two public datasets have demonstrated the effectiveness of our approach. Our GarageWorld, along with the reconstructed 3D model and real-time rendering, enables a set of applications, including training data generation and testbed for autonomous driving algorithms, real-time assistance for autonomous vehicle localization, navigation, and parking, as well as VFX production. Our current contributions mainly focus on the perception of the world, and it enables downstream recognition tasks. In the future, we will also explore garage generation, keeping pushing the boundary of garage modeling and accomplishing a closure from perception, recognition and generation.

### v2

@@ 7 Limitations and Discussions
Although our rendering results exhibit commendable realism and we showcase various applications, our workflow still possesses certain limitations. Here we present a detailed analysis and explore further potential applications.
Firstly, our method focuses on large-scale garage scenes and relies on our lightweight 3D scanner to provide high-quality LiDAR point clouds, color images, and corresponding camera information. Although we have also validated the effectiveness of our method on other open-source datasets, it is important to explore the use of even lighter devices for scanning and rendering large-scale scenes, such as smartphones equipped with depth sensors. Additionally, we have currently collected data from eight large-scale garages, but it is necessary to gather more garage data to contribute to the community and facilitate further research on garage modeling and rendering.
As a method based on image rendering, our approach achieves highly realistic rendering effects, almost indistinguishable from real scenes. However, the existing pipeline does not support modifying lighting conditions. This requires us to carefully design the shooting process according to the lighting conditions, limiting the applicability of our method in various scenarios. Enabling rendering and editing of large-scale scenes under different lighting conditions is a meaningful direction that deserves further investigation. Our open-source dataset provides a foundation for the community to conduct research in these directions, allowing for advancements in this field.
Furthermore, there are additional directions and applications worth exploring. Inspired by SMERF [17], one potential research direction is to investigate streaming transmission methods for Gaussian kernels, enabling the distribution and on-the-fly rendering of large-scale scene data. Additionally, there are city generation methods such as InfiniCity [45] and CityDreamer [85], which leverage Generative Adversarial Networks (GANs) [26] to achieve rapid modeling of large-scale scenes. We intend to study large-scale scene generation based on 3D Gaussian representations. This fully explicit representation can be easily integrated into existing computer graphics workflows and achieve superior rendering effects.
@@ 8 Conclusion
For many of us, our daily lives begin with a safe departure from a garage and end with a safe arrival. The garage serves as the origin of our journey to innovation. This paper has contributed a handheld Polar device for data collection, a GarageWorld dataset, LiDAR-assisted Gaussian splatting for scene representation, and an LOD-based rendering technique that allows web-based rendering on consumer-level devices. Benefiting from these innovations, we successfully reconstruct various garages with diverse and challenging environments, allowing real-time rendering from any viewpoint on lightweight devices. Experimental results on the collected and two public datasets have demonstrated the effectiveness of our approach. Our GarageWorld, along with the reconstructed 3D model and real-time rendering, enables a set of applications, including training data generation and testbed for autonomous driving algorithms, real-time assistance for autonomous vehicle localization, navigation, and parking, as well as VFX production. Our current contributions mainly focus on the perception of the world, and it enables downstream recognition tasks. In the future, we will also explore garage generation, continuing to push the boundary of garage modeling and accomplishing a closure from perception, recognition, and generation.
Appendix A Specifications and Calibration of Polar Device
Below, we provide the configuration and calibration details of our Polar scanner. The data acquisition unit comprises an Ouster OS0-128 REV6 LiDAR, an Insta360 ONE RS 1-inch 360 Edition Camera, and an Xsens MTi-630 IMU. Time synchronization across sensors is achieved using DEITY TC-1 timecode generators and simulated analog GPS NMEA time signals. The data collection unit and processing unit weigh 1250g and 850g, respectively, excluding the battery. The total apparatus weight does not exceed 3.2 kg, making it ideal for both handheld and wearable scanning applications. Additionally, we have engineered an over-shoulder support system to enhance operator stability during scans.
For the intrinsic parameter calibration of the fisheye camera and the IMU sensor, we leverage the program provided by OpenCV and the Allan Variance ROS toolbox [59, 22, 21, 50, 56], respectively. The extrinsic calibration between the fisheye camera and the IMU uses the Kalibr calibration program [59]. Due to the sparsity, noisiness, and uncolored nature of the point cloud data collected by our LiDAR sensor, it is hard to establish correspondences between the LiDAR data and the images captured by the fisheye camera, making relative pose estimation between the sensors difficult. To address this issue, we introduce an additional sensor, a FARO laser scanner. This scanner provides dense, accurate, and colored point clouds and thus functions as a bridge for the relative pose estimation between the LiDAR sensor and the fisheye camera. Specifically, we adopt a checkerboard as our calibration scene and capture data from it using the FARO laser scanner and Polar device. The relative pose between the FARO laser scanner and the LiDAR sensor in our Polar device is achieved by point cloud registration, and that between the FARO laser scanner and the fisheye camera is obtained by establishing color correspondences between the point cloud captured by the FARO sensor and the fisheye image. By connecting the transformation from FARO to LiDAR and from FARO to the fisheye camera, we can finally calculate the extrinsic parameters between the LiDAR and the fisheye camera. We repeat this process twelve times in our implementation to reduce experimental error. It is important to note that FARO was only used to calibrate between the fisheye camera and the LiDAR before we start scanning a garage, and the process only needs to be conducted once.
Appendix B Data Preprocessing and Mesh cleanup
To ensure data privacy, we anonymized vehicle license plate information recorded in the underground garage data collection. Specifically, we used image blurring to hide license plate numbers from identification. This step is essential to protect individual privacy and comply with data protection regulations. To remove the dynamic objects in the collected data, we apply Segment Anything [37] to the images. For the point cloud data, we estimate whether the points scanned at one LiDAR frame are also observed at other frames, similar to Schauer and Nüchter [63]. For mesh cleaning, we construct a kd-tree from the point cloud of LIV-SLAM and calculate the nearest neighbour distance of each mesh face (k = 1). If the distance exceeds 0.1m, we remove the face.
Appendix C Details of depth calculation
To accurately calculate depth, we treat each Gaussian as a probability density function. Following [100], we get the depth of each Gaussian primitive, $d_{i}$, by calculating the expected value of depth under this probability density as:
[…]
where $g_{i}$ is the Gaussian function, $\boldsymbol{x}=[x_{0},x_{1},x_{2}]^{T}$ is the coordinate of the space in the ray space, where $x_{0},x_{1}$ are the pixel coordinates, $x_{2}$ is oriented in the same direction as the pixel ray. As in [100], $t=x_{2}/l$ is the depth of $\boldsymbol{x}$ in the camera coordinate system and $l=\sqrt{x_{0}^{2}+x_{1}^{2}+1}$. We denote the center of the Gaussian in ray space as $\boldsymbol{p}=[p_{0},p_{1},p_{2}]$ and the covariance matrix as $\boldsymbol{\boldsymbol{\Sigma}}$, by defining $\boldsymbol{y}=\boldsymbol{x}-\boldsymbol{p}$, the Eqn. 9 can be simplified as:
[…]
Then, $g(\boldsymbol{y}+\boldsymbol{p})$ can be expanded as:
$g(\boldsymbol{y}+\boldsymbol{p})=e^{-Ay_{2}^{2}+By_{2}+C}=e^{C}\cdot e^{-Ay_{2}^{2}+By_{2}},$ (11)
where
$\begin{split}A&=\frac{1}{2}\left(\boldsymbol{\Sigma}^{-1}\right)_{2,2}\\ B&=-\left(\boldsymbol{\Sigma}^{-1}\right)_{2,0}y_{0}-\left(\boldsymbol{\Sigma}^{-1}\right)_{2,1}y_{1}\\ C&=-\frac{1}{2}\left(\boldsymbol{\Sigma}^{-1}\right)_{0,0}y_{0}^{2}-\frac{1}{2}\left(\boldsymbol{\Sigma}^{-1}\right)_{1,1}y_{1}^{2}-\left(\boldsymbol{\Sigma}^{-1}\right)_{1,0}y_{0}y_{1}\\ .\end{split}$ (12)
By integrating an arbitrary Gaussian function, we obtain:
[…]
Next, substituting Eqn. 13 and Eqn. 14 into Eqn. 10 enables us to obtain:
[…]
Finally, through simplification, Eqn. 15 can be transformed into Eqn. 4.
Before depth supervision, we normalize both the captured depth and predicted depth to $0-1$ as follows:
$R(D)=\begin{cases}\frac{1}{2\beta}D&(D<\beta)\\ 1-\frac{\beta}{2D}&(D\geq\beta)\end{cases},$ (16)
where $\beta=10$, as the depth values within 10 meters are accurate. In this way, the same depth error results in a greater loss when closer to the camera than at further distances. It effectively avoids excessive depth loss for distant Gaussians due to inherently large depth values. This normalization also facilitates the storage of depth maps.